A 85 Kg block of ice is being pushed in a straight rough road. The force exerted on the ice 300 N and acceleration of the entire system is 10 m/s^2. What is the friction force between the ice and the rough road.


1
Expert's answer
2022-05-30T15:17:36-0400
"ma=F-F_f"

"F_f=F-ma\\
=\rm 300\: N-8.5\: kg\times 10\: m/s^2=215\:\rm N"